16 years from now pia's age will be twice her age 12 years ago whats her current age

Answer by TimothyLamb(4379)

x = pia's age now
x + 16 = 2(x - 12)
x + 16 = 2x - 24
x = 40
x = pia's age now = 40
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php